A thin metallic spherical shell of radius 0.387 m has a total charge of 9.03 × 10-6 C placed on it. At the center of the shell is placed a point charge of 1.65 × 10-6 C. What is the electric field at a distance of 0.995 m from the center of the spherical shell?Express your answer in N/C
Answer #1

here total charge Q = Qin + qo

Q = 9.03 *10^-6 + 1.65 *10^-6 C

Q = 10.68 *10^-6 C


Electric field due point charge is given by E = QK/r^2


where k is permittivity constant = 9e9

so

E = (9*10^9 * 10.68 *10^-6)/(0.995* 0.995)

E = 97088.45 N/C is the electric field at a point of 0.995 m